Blog

MOVE Software Update

 

 

Email from Larry -  

So I went out to the mesa today and put in those changes to the code I had mentioned.   This version (which _appears_ to work OK) has now replaced "newmove.exe".

 

I have also made this the default move.  I'm hoping the changes will fix the problem and I really can't wait months to get this problem debugged.  If things go belly-up, Han, Avner or Croll can run oldmove.exe instead -- it's still there and they are the only ones on for the rest of the month.

 

Question -- If I exit MOVE and run OLDMOVE, do I need to restart LOIS to get the communication going?  Or can I do a tel_id to LOIS.  That used to work at one time.

 

Larry

 

Minor Maintenance

Added Helium to the Mimir Compressor at 9:40am

Starting cooling the primary mirror this morning as well. Will check for leaks before I leave for the hill.

I have completed the migration of documentation from Lux to Jumar. Hopefully I got everything. - BWT

Jim cleaned the mirror today.

 

Brand new Gas Cylinder from PraxAir, Cylinder weight 197.0lbs, 620 PSI

After cleaning

183lbs and 600 PSI

 

- Mirror Cleaning with CO2

Jim cleaned the mirror today.

Cylinder Weight - 173 lbs before cleaning. This is with cart and cap, but no regulator.

Jim and I discusses safety issues regarding and resolved to do the following

  1. Tub Fans will be on.
  2. Tub Exhaust Fan will be on
  3. Jim will make a new lanyard to protect the mirror if the nozzle is dropped for some reason.
  4. Jim will make a new lanyard to protect the operator if the hose comes loose from the nozzle.

Cylinder Weight - 154 lbs after cleaning.

Instrument Maintenance 09/02/2014

Mimir Cooling Start

    • Turned on BA Gauge
    • Pegged the detector temperature to the current temperature, make sure to turn the heater on.
    • Set the temperature set point for the second stage to 10.0K
    • Changed the temper log file data to reflect the current date i.e. 20140902.tc.log in this case.
    • Changed plottempa.pro to look at new temper file.
    • Rebuilt the IDL virtual machine program to push the web pages.
    • Moved the Camera, Slit, and Decker to home
    • Started the runstage program to move the mechanisms while cooling.
    • Started the compressor at 13:13 09/02/2014.
    • Changed the Amperage for the BA gauge to 6.5 mA at 14:15.
    • Fix the crontab so the BA gauge would automatically turn on and off.
    • 16:09, started the detector to ramp at 0.1 K/m to 33.5K
    • 16:36, Valved off the vacuum pumps

 

PRISM Maintenance

    • Replaced the fans on the instrument mechanisms.
    • Increased the tension on FW3 detent arm.(Thanks Dan!)
Old Blog Post from Lux
Last changed: Dec 09, 2013 11:44 by Brian W. Taylor

I added a new section on what to do if for Mimir. The first note is how to clean up the MimirInst daemon that can be found here

http://lux.lowell.edu:8090/display/PerkinsInfo/Mimir+Instrument+Hang

Posted at 09 Dec @ 11:42 AM by Brian W. Taylor | 0 comments | Edit
  2013/11/26
Last changed: Nov 26, 2013 11:09 by Brian W. Taylor

Added Rich's 2004 Control System Manual to the Confluence pages. It can be downloaded with this link Perkins_Manual.pdf

Posted at 26 Nov @ 11:09 AM by Brian W. Taylor | 0 comments | Edit
  2013/03/18
Last changed: Mar 18, 2013 11:47 by Brian W. Taylor

Mimir has been valved off of the vacuum pumps and the detector has been put into a .1K/min ramp to 33.5.

Posted at 18 Mar @ 11:46 AM by Brian W. Taylor | 0 comments | Edit
Last changed: Mar 18, 2013 10:23 by Brian W. Taylor

The BA gauge has gone into the Error state, E03, which means that the gauge has saturated. Will bump up the current and see if it will read. Otherwise it will be time to take Mimir off the pumps.

Posted at 18 Mar @ 10:21 AM by Brian W. Taylor | 0 comments | Edit
Last changed: Mar 18, 2013 09:18 by Brian W. Taylor

Mimir started cooling at 9:12am 03/18/2013

Posted at 18 Mar @ 9:18 AM by Brian W. Taylor | 0 comments | Edit
  2012/08/17
Last changed: Aug 17, 2012 23:22 by Brian W. Taylor
Labels: mimir, aeroquip, return, cold, head, cooling, test

Did a cold-head purge tonight and to clean out the head. In the process of  attaching the return line to the cold head, the aeroquip connector failed. The caused a loss of most of the Helium in the system. Got this fixed and purged both the line that was the cause, the coldhead, and all of the lines. Once this was done I started a cool down test to see if things where operating correctly. 

During the initial coldhead purge the throughput for the compressor was 22 SCFM 

Start of the test

20120817 21:59:13 Setpoint: 305.00 230.00, Ramp: 1 1, Rate: 0.10 0.10, R Status: 0 0 HeaterA: 6.00% of range 3  HeaterB: Analog Out 0.00
20120817 21:59:13 A0 305.00K B0 301.74K C1 303.61K C2 475.00K C3 128.44K C4 293.16K C5 305.07K C6 302.74K C7 179.27K C8 304.05K

1st stage starting a slow down

20120817 22:34:13 Setpoint: 305.00 230.00, Ramp: 1 1, Rate: 0.10 0.10, R Status: 0 0 HeaterA: 0.00% of range 0  HeaterB: Analog Out 0.00
20120817 22:34:13 A0 189.53K B0 301.70K C1 303.44K C2 475.00K C3 128.44K C4 292.82K C5 303.57K C6 302.72K C7 179.06K C8 232.61K
20120817 22:34:13 TC1 1.00e-03T TC2 0.00e+00T CG 7.40e-09T

Powering down,  things look good.

20120817 23:18:13 Setpoint: 305.00 230.00, Ramp: 1 1, Rate: 0.10 0.10, R Status: 0 0 HeaterA: 0.00% of range 0  HeaterB: Analog Out 0.00
20120817 23:18:13 A0 109.29K B0 300.58K C1 301.93K C2 475.00K C3 128.44K C4 292.46K C5 299.07K C6 302.48K C7 178.88K C8 222.73K
20120817 23:18:13 TC1 1.00e-03T TC2 0.00e+00T CG 8.64e-10T

Posted at 17 Aug @ 10:38 PM by Brian W. Taylor | 0 comments | Edit
  2012/08/14
Last changed: Aug 14, 2012 18:16 by Brian W. Taylor
Labels: mimir, cold-head, adsorber

Performed the following Maintenance on Mimir, with help from Steve Winchester

1) Pulled the cold-head and rebuilt parts of it. Replaced the Delrin bushings for both upper and lower yolk holds. Most of the large o-rings and the displacer seal.

2) Replaced the adsorber.

3) Flushed the Helium lines from the compressor

Next step will be to flush the cold-head and then do a test of the system before starting the cool down.

Posted at 14 Aug @ 6:12 PM by Brian W. Taylor | 0 comments | Edit
  2012/07/17
Last changed: Jul 17, 2012 14:51 by Brian W. Taylor

Started Mimir warming today.

The Current Plan

  1. run the detector up so it is above some of the optics
  2. disconnect the helium hoses
  3. turn off the compressor

I will leave the pumps on until it gets past some of the major
volatiles, then turn them off. Since we will not be cracking
the vacuum will continue to pump occasionally until about 18th
of Aug. Do a pump and bake from the 18th til the 25th then start
cooling to plan to be cold before Sept 3rd.

Posted at 17 Jul @ 2:47 PM by Brian W. Taylor | 0 comments | Edit
  2012/06/23
Last changed: Jun 23, 2012 08:59 by Brian W. Taylor

New Electronics mount installed the 06/14/2012. 

Design Problems:

1) Need to shave 0.25" off the cryostat interface of the bottom brackets

2) Need to increase the hole size to a drill size "T"

Once those two things were fix the mount went on very nicely and everything worked
well except for the one amplifier line on Q1OUT2- or pin W of the Red cable this supplies
the upper left hand quadrant, Column 2. After some work, I was able to get the quadrant
to work and latched down the cables. That flaky connection really did not like the new
orientation.

Mimir installed on Perkins 06/22/2012

Ralph came out and balanced Mimir with the new electronics mount configuration. 

However when Mimir was installed onto the telescope there was significant pattern
noise that was new to the array. Tracking it down, it was discovered that the bottom
brackets of the electronics mount need some type of insulation between the top of
the bracket, that had been shaved off, and the cryostat.

During the troubleshooting the wire finally broke for good and no amount of
fiddling would get the connection back to working. I pulled the cables and stripped
the heat shrink. At this point Dan had arrived and volunteered to solder the
wire back in place. This was done and everything was clean up and Dan was
able to observer with no issues.

Posted at 23 Jun @ 8:59 AM by Brian W. Taylor | 0 comments | Edit
  2012/04/02
Last changed: Apr 02, 2012 17:59 by Brian W. Taylor

 I have installed a BlueRay RW drive at the Perkins. Storage
capacity will be 25GB for a single layer disk and 50GB for a double
layer disk. See the attached for the formats and capacities accepted
for the drive.

BlueRay_DriveFormats.pdf

Posted at 02 Apr @ 5:59 PM by Brian W. Taylor | 0 comments | Edit
  2012/03/27
Last changed: Mar 27, 2012 11:13 by Brian W. Taylor

I have removed the GRISMs from the lower GRISM wheel and have installed the SLOAN Filters

  • Position 1 - Empty
  • Position 2 - SDSS u
  • Position 3 - SDSS g
  • Position 4 - SDSS r
  • Position 5 - SDSS i
  • Position 6 - SDSS z

The numbers below were taken during the day.

Dome Flat Times

   SDSS u - Narrow Band Lamps at 70%  30

Normal Broad Band Lamps

   SDSS g - 15s ~22000 ADU

   SDSS r - 10s ~21000 ADU

   SDSS i - 12s ~21000 ADU

   SDSS z - 20s ~20000 ADU

  

Posted at 27 Mar @ 11:13 AM by Brian W. Taylor | 0 comments | Edit
  2011/03/14
Last changed: Mar 14, 2011 14:24 by Brian W. Taylor

Added Helium to Mimir to boost the pressure to 60-70 psi on return

Posted at 14 Mar @ 2:23 PM by Brian W. Taylor | 0 comments | Edit
  2011/01/19
Last changed: Jan 19, 2011 17:29 by Brian W. Taylor

1/19/11 10:02 AM - Problem with the K-HWP is due to a mechanical interference.
Motion at the Flat screen position seems to work well. But not many other places.
The failure mode is that when it fails the HWP kicks the filter. This changes the encoder
position. From 13 to 12
A different mode is that when the HWP found home the encode changed from 13 to 15

 My Note to Dan regarding the failure.

January 19, 2011 11:30:33 AM MST
Hi Dan,
Here is what is happening. 
The filter wheels move into the correct position and the encoder
bit show that things are correct. When you try to home the HWP it
will fail and this does return an error. The interference causes the
P wheel to kick which is reported in the encoder for the P wheel. It 
will either switch to 12 or 15 instead of 13. 
The real issue is when the HWP finds home be causes the
kick in a move after the home move. There is no way to tell a 
failure unless you check the filter encoders to see if the wheel
has moved.

I have had very few places where I can get it to work. It
seems as if the flat screen is the best location.

Brian

Posted at 19 Jan @ 5:25 PM by Brian W. Taylor | 0 comments | Edit
  2011/01/07
Last changed: Jan 07, 2011 14:07 by Brian W. Taylor

Added Helium to the compressor. Bump from about 30 PSI to 100 PSI on the return.

Posted at 07 Jan @ 2:06 PM by Brian W. Taylor | 0 comments | Edit
  2010/12/18
Last changed: Dec 18, 2010 17:04 by Brian W. Taylor

The compressor failed this afternoon at 3:40pm, I was able to get restarted before 5 pm so the detector did not heat up too much. The failure was in a mode that the remote operated circuit breaker did not restart the compressor.

This is a power module deep in the electronics cage that required a button push to restart the compressor. This is only the second time that this has been required but I think that it calls for a query to Oxford instruments to find out what is going on with it.

Posted at 18 Dec @ 5:01 PM by Brian W. Taylor | 0 comments | Edit